Calculating Interest Rates: Solve for the unknown interest rate in each of the following:

Present Value Years Interest Rate Future Value
              217 3                 293
              432 10             1,053
        41,000 16         162,181
        54,382 19         483,500

Calculating Future Values: Compute the future value of $3550 compounded annually for

a. 10 years at 6%

b. 10 years at 8%

c. 20 years at 6%

d. Why is the interest earned in part (c) not twice the amount earned in part (a)?
Answer #1
As per HOMEWORKLIB RULES we answer one question per post. Kindly post remaining questions in next post
Dear Student
Thank you for using homeworklib
Please find below the answer
Statementshowing Computations
Paticulars
a)
293 = 217(1+r/100)^3
1.3502 = (100+r/100)^3
1.1053 = (100+r)/100
110.53 = 100 + r
r = 10.53%
b)
432 = 1053(1+r/100)^10
2.4375 = (1+r/100)^10
1.0932 = (100 +r)/100
109.32 - 100 =r
r = 9.32%
c)
162,181 = 41000(1+r/100)^16
3.9556 = (1+r/100)^16
1.0897 = (100 +r)/100
108.97 = 100 + r
r = 8.97%
d)
483500 = 54382(1+r/100)^19
8.8908 = (100 +r/100)^19
1.1219 = (100 +r)/100
112.19 - 100 =r
r = 12.19%